# Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization

### v2

@@ B.2.1 Baselines
Google OR-Tools (Furnon & Perron, 2024) is an open-source software suite designed to address various combinatorial optimization problems. This toolkit offers a comprehensive selection of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges. Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy. For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents. We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization. The solving time was set as $\{30,60,300,600\}$ seconds for $N=\{50,100,500,1000\}$, respectively.
@@
The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network-integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes. This approach helps the deep reinforcement learning model to make informed node selections during route planning. We adapt the original model to handle OMDCPDP with multiple agents akin to the sequential planning of Son et al. (2024).
@@
The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) introduces a cooperative PDP with multiple vehicle agents. This framework is trained via a centralized (MA)RL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes. We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison. For the decoder and training phases, we kept the same random conflict handler, and we retained the hyperparameters detailed in the original study overall.
@@ B.2.2 Datasets
Neural baselines were trained with the specific number of nodes $N$ and number of agents ${M}$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(5,20)$. As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. Similarly, we sample ${M}$ initial vehicle locations from the same distribution. We set the demand $d_{i}$ for each customer to $1$ and the capacity of each vehicle to $3$. This emulates realistic settings in which a single package per customer will be picked up and delivered.
@@
Testing is performed on $1000$ new instances for each setting of in-distribution $N$ and ${M}$ in Table 1 with the distributions from the training settings. For large-scale generalization in Table 2, we generate $100$ new instances.
@@ B.2.3 PARCO Network Hyperparameters
Most hyperparameters are kept similar to Section B.1.3.
@@
Initial Embedding. This layer projects initial raw features to hidden space. For depots, the initial embeddings encode the location $o_{m}$ and the respective vehicle’s capacity $Q_{m}$. For pickup nodes, the initial embeddings encode the location and paired delivery nodes’ location. For delivery nodes, the initial embeddings encode the location and paired pickup nodes’ location. Main Encoder. We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP. Note that unlike the encoder layer, which acts between all $M+N$ tokens, communication layers are lighter because they communicate between ${M}$ agents.
@@
We employ the Priority-based Conflict Handler guided by the model output probability for managing conflicts with priority given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
@@ B.2.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $10^{5}$ randomly generated data. Training takes less than 5 hours in our configuration.
@@ B.3.1 Baselines
We implement the mathematical model described above in the exact solver Gurobi (Gurobi Optimization, LLC, 2024) with a time budget of 60 and 600 seconds per instance. However, with both time budgets, Gurobi is only capable of generating solutions to the FFSP20 instances, similar to the findings made by Kwon et al. (2021) for the CPLEX solver.
@@
The Random and Shortest Job First (SJF) heuristics are simple construction strategies that build valid schedules in an iterative manner. Starting from an empty schedule, the Random construction heuristic iterates through time steps $t=0,\dots T$ and stages $i=1\dots S$ and randomly assigns jobs available at the given time to an idle machine of the respective stage until all jobs are scheduled. Likewise, the SJF proceeds by assigning job-machine pairs with the shortest processing time first. [^2] To obtain results for the heuristics and metaheuristics, we used the implementation of Kwon et al. (2021), provided in the official GitHub repository of the paper: https://github.com/yd-kwon/MatNet
@@
Genetic Algorithms are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008). The GA iteratively improves multiple candidate solutions called chromosomes. Each chromosome consists of $S\times N$ real numbers, where $S$ is the number of stages and $N$ is the number of jobs. For each job at each stage, the integer part of the corresponding number indicates the assigned machine index, while the fractional part determines job priority when multiple jobs are available simultaneously. Child chromosomes are created through crossover, inheriting integer and fractional parts independently from two parents. Mutations, applied with a 30% chance, use one of four randomly selected methods: exchange, inverse, insert, or change. The implementation uses 25 chromosomes. One initial chromosome is set to the Shortest Job First (SJF) heuristic solution and the best-performing chromosome is preserved across iterations. Each instance runs for 1,000 iterations.
@@
Finally, Particle Swarm Optimization iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh & Mahapatra, 2012). In this implementation, particles use the same representation as GA chromosomes. The algorithm employs 25 particles, with an inertial weight of 0.7 and cognitive and social constants set to 1.5. One initial particle represents the SJF heuristic solution. Like GA, PSO runs for 1,000 iterations per instance.
@@
We benchmark PARCO mainly against MatNet (Kwon et al., 2021), a state-of-the-art NCO architecture for the FFSP. MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2019a). It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable of encoding nodes of different types in bipartite graph-like machines and jobs in the FFSP. Kwon et al. (2021) train MatNet using POMO (Kwon et al., 2020).
@@ B.3.2 Datasets
We follow the instance generation scheme outlined in Kwon et al. (2021) sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$. For the first three FFSP instance types shown in Table 1 we also use the same instance sizes as Kwon et al. (2021) with $N=20,50\text{ and }100$ jobs and $M=12$ machines which are spread evenly over $S=3$ stages. To test for agent sensitivity in the FFSP, we fix the number of jobs to $N=50$ but alter the number of agents for the last three instance types shown in Table 1. Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $M_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
@@ B.3.3 PARCO Network Hyperparameters
To solve the FFSP with our PARCO method, we use a similar encoder as Kwon et al. (2021). The MatNet encoder generates embeddings for all machines of all stages and the jobs they need to process, plus an additional dummy job embedding, which can be selected by any machine in each decoding step to skip to the next step. To compare PARCO with MatNet, we use similar hyperparameters for both models. We use $L=3$ encoder layers, generating embeddings of dimensionality $d_{h}=256$, which are split over $h=16$ attention heads in the MHA layers. Further, we employ Instance Normalization (Ulyanov et al., 2016) and a feed-forward network with 512 neurons in the transformer blocks of the encoder.
@@
The machines are regarded as the agents in our PARCO framework. As such, their embeddings are used as queries $\bm{q}$ in the Multiple Pointer Mechanism Equation 7, while job embeddings are used as the keys and values. In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle. Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Equation 6.
@@ B.3.4 PARCO Training Hyperparameters
Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Equation 9. We use the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme. We train separate models for the environment configurations used in Table 1. We train models corresponding to environments with 20 jobs for 100, with 50 jobs for 150 and with 100 jobs for 200 epochs. In each epoch, we train the models using 1,000 randomly generated instances split into batches of size 50.333Note: to avoid OOMs, For FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated.
@@
The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO. As one can see in Figure 5(a), MatNet requires a decoder forward pass for each machine to schedule a job on each of them. In contrast, as detailed in Figure 5(b), PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
@@ B.4.2 Software
We used Python 3.12, PyTorch 2.5 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a). The operating system is Ubuntu 24.04 LTS.
@@ B.5 Source Code
B.5.1 Licenses
We provide the summary of licenses Table 3.
@@ B.5.1 Licenses
OR-Tools (Furnon & Perron, 2024) Code Apache License, Version 2.0
AM (Kool et al., 2019a) Code MIT License
ET (Son et al., 2024) Code Available on Github
DPN (Zheng et al., 2024a) Code MIT License
DRLLi (Li et al., 2022) Code Available on Github
2D-Ptr (Liu et al., 2024b) Code/Dataset Available on Github
HAM (Li et al., 2021a) Code MIT License
Gurobi (Gurobi Optimization, LLC, 2024) Code Commercial license (free for academic use)
MatNet (Kwon et al., 2021) Code/Dataset MIT License
RL4CO (Berto et al., 2024a) Library MIT License
B.5.2 Code Release
We value open reproducibility and release our source code at https://github.com/ai4co/parco.

### v3

@@ B.2.1 Baselines
Google OR-Tools [19] is an open-source software suite designed to address various combinatorial optimization problems. This toolkit offers a comprehensive selection of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges. Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy. For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents. We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization. The solving time was set as $\{30,60,300,600\}$ seconds for $N=\{50,100,500,1000\}$, respectively.
@@
The Heterogeneous Attention Model (HAM) [54] utilizes a neural network-integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes. This approach helps the deep reinforcement learning model to make informed node selections during route planning. We adapt the original model to handle OMDCPDP with multiple agents akin to the sequential planning of Son et al. [84].
@@
The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) [108] introduces a cooperative PDP with multiple vehicle agents. This framework is trained via a centralized (MA)RL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes. We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison. For the decoder and training phases, we kept the same random conflict handler, and we retained the hyperparameters detailed in the original study overall.
@@ B.2.2 Datasets
Neural baselines were trained with the specific number of nodes $N$ and number of agents M{M} they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(5,20)$. As we show in the Table˜1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. Similarly, we sample M{M} initial vehicle locations from the same distribution. We set the demand $d_{i}$ for each customer to $1$ and the capacity of each vehicle to $3$. This emulates realistic settings in which a single package per customer will be picked up and delivered.
@@
Testing is performed on $1000$ new instances for each setting of in-distribution NN and M{M} in Table˜1 with the distributions from the training settings. For large-scale generalization in Table˜2, we generate $100$ new instances.
@@ B.2.3 PARCO Network Hyperparameters
Most hyperparameters are kept similar to §˜B.1.3.
@@
Initial Embedding. This layer projects initial raw features to hidden space. For depots, the initial embeddings encode the location $o_{m}$ and the respective vehicle’s capacity $Q_{m}$. For pickup nodes, the initial embeddings encode the location and paired delivery nodes’ location. For delivery nodes, the initial embeddings encode the location and paired pickup nodes’ location. Main Encoder. We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP. Note that unlike the encoder layer, which acts between all $M+N$ tokens, communication layers are lighter because they communicate between M{M} agents.
@@
We employ the Priority-based Conflict Handler guided by the model output probability for managing conflicts with priority given to the agent whose probability of selecting the conflicting action is the highest (see §˜4.6).
@@ B.2.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO [43] with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer [46] with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $10^{5}$ randomly generated data. Training takes less than 5 hours in our configuration.
@@ B.3.1 Baselines
We implement the mathematical model described above in the exact solver Gurobi [24] with a time budget of 60 and 600 seconds per instance. However, with both time budgets, Gurobi is only capable of generating solutions to the FFSP20 instances, similar to the findings made by Kwon et al. [50] for the CPLEX solver.
@@
The Random and Shortest Job First (SJF) heuristics are simple construction strategies that build valid schedules in an iterative manner. Starting from an empty schedule, the Random construction heuristic iterates through time steps $t=0,\dots T$ and stages $i=1\dots S$ and randomly assigns jobs available at the given time to an idle machine of the respective stage until all jobs are scheduled. Likewise, the SJF proceeds by assigning job-machine pairs with the shortest processing time first. [^2] To obtain results for the heuristics and metaheuristics, we used the implementation of Kwon et al. [50], provided in the official GitHub repository of the paper: https://github.com/yd-kwon/MatNet
@@
Genetic Algorithms are metaheuristics widely used by the OR community to tackle the FFSP [38]. The GA iteratively improves multiple candidate solutions called chromosomes. Each chromosome consists of $S\times N$ real numbers, where $S$ is the number of stages and $N$ is the number of jobs. For each job at each stage, the integer part of the corresponding number indicates the assigned machine index, while the fractional part determines job priority when multiple jobs are available simultaneously. Child chromosomes are created through crossover, inheriting integer and fractional parts independently from two parents. Mutations, applied with a 30% chance, use one of four randomly selected methods: exchange, inverse, insert, or change. The implementation uses 25 chromosomes. One initial chromosome is set to the Shortest Job First (SJF) heuristic solution and the best-performing chromosome is preserved across iterations. Each instance runs for 1,000 iterations.
@@
Finally, Particle Swarm Optimization iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration [81]. In this implementation, particles use the same representation as GA chromosomes. The algorithm employs 25 particles, with an inertial weight of 0.7 and cognitive and social constants set to 1.5. One initial particle represents the SJF heuristic solution. Like GA, PSO runs for 1,000 iterations per instance.
@@
We benchmark PARCO mainly against MatNet [50], a state-of-the-art NCO architecture for the FFSP. MatNet is an encoder-decoder architecture, which is inspired by the attention model [47]. It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable of encoding nodes of different types in bipartite graph-like machines and jobs in the FFSP. Kwon et al. [50] train MatNet using POMO [49].
@@ B.3.2 Datasets
We follow the instance generation scheme outlined in Kwon et al. [50] sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$. For the first three FFSP instance types shown in Table˜1 we also use the same instance sizes as Kwon et al. [50] with $N=20,50\text{ and }100$ jobs and $M=12$ machines which are spread evenly over $S=3$ stages. To test for agent sensitivity in the FFSP, we fix the number of jobs to $N=50$ but alter the number of agents for the last three instance types shown in Table˜1. Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $M_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
@@ B.3.3 PARCO Network Hyperparameters
To solve the FFSP with our PARCO method, we use a similar encoder as Kwon et al. [50]. The MatNet encoder generates embeddings for all machines of all stages and the jobs they need to process, plus an additional dummy job embedding, which can be selected by any machine in each decoding step to skip to the next step. To compare PARCO with MatNet, we use similar hyperparameters for both models. We use $L=3$ encoder layers, generating embeddings of dimensionality $d_{h}=256$, which are split over $h=16$ attention heads in the MHA layers. Further, we employ Instance Normalization [87] and a feed-forward network with 512 neurons in the transformer blocks of the encoder.
@@
The machines are regarded as the agents in our PARCO framework. As such, their embeddings are used as queries $\bm{q}$ in the Multiple Pointer Mechanism Eq.˜6, while job embeddings are used as the keys and values. In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle. Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Eq.˜5.
@@ B.3.4 PARCO Training Hyperparameters
Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Eq.˜8. We use the Adam optimizer [46] with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme. We train separate models for the environment configurations used in Table˜1. We train models corresponding to environments with 20 jobs for 100, with 50 jobs for 150 and with 100 jobs for 200 epochs. In each epoch, we train the models using 1,000 randomly generated instances split into batches of size 50.333Note: to avoid OOMs, for FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated.
@@
The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO. As one can see in Fig.˜6(a), MatNet requires a decoder forward pass for each machine to schedule a job on each of them. In contrast, as detailed in Fig.˜6(b), PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
@@ B.4.2 Software
We used Python 3.12, PyTorch 2.5 [78] coupled with PyTorch Lightning [17] with most code based on the RL4CO library [5]. The operating system is Ubuntu 24.04 LTS.
@@ B.5 Source Code
We value open reproducibility and release our source code at https://github.com/ai4co/parco. We also provide the summary of licenses in Table˜3.
@@
OR-Tools [19] Code Apache License, Version 2.0
AM [47] Code MIT License
ET [84] Code Available on Github
DPN [104] Code MIT License
DRLLi [55] Code Available on Github
2D-Ptr [63] Code/Dataset Available on Github
HAM [54] Code MIT License
Gurobi [24] Code Commercial license (free for academic use)
MatNet [50] Code/Dataset MIT License
RL4CO [5] Library MIT License
Appendix C Additional Materials
C.1 Further Discussions
Follow-ups
We acknowledge there are works based on or inspired by PARCO already out in the wild at the time of publication. These include extending to profiled VRP [29, 30], mixed-shelves picker-routing [67], and self-improvement for better performance [68]. We hope to see even more in the future from diverse research groups, and we remain available for discussions, including through the AI4CO Slack.
Possible future directions
In this paragraph, we will share a short list of possible additional future directions for PARCO. One promising direction for NCO is to employ Large Language Models (LLMs) for automating the design of algorithms and particularly heuristics [62, 97, 27]. Several components of PARCO could be designed by LLMs, including the conflict handling mechanism or attention biases dependent on specific problems [31, 90] as done by Tran et al. [86]. Of interest would also be extensions to other problem variants [99]. Integrating parallel decoding into search methods, as NDS [28], may also yield much more efficient models and enable them to capture different agents.
C.2 Comparison with Decentralized and Graph-Based Communication Methods
We additionally compare PARCO with decentralized methods and alternative communication models like Graph Neural Networks (GNNs). We experiment on the min-max traveling salesman problem (mTSP), benchmarking PARCO against notable decentralized methods like GNN-DisPN [16], DAN [10] and GNN-based communication models like ScheduleNet [75].
The results, shown in Table˜4 with instances from Park et al. [75] as well as in the mTSPLib [^4] https://profs.info.uaic.ro/mihaela.breaban/mtsplib/MinMaxMTSP/ in Table˜5 – additionally adding the HGA solver for further comparison from Mahmoudinazlou and Kwon [72] – demonstrate that PARCO consistently outperforms these methods in solution quality. Centralized training with parallel decoding, as used in PARCO, offers distinct advantages by enabling global coordination and highly efficient solution construction. For example, GNN-based communication in models like ScheduleNet incurs significantly higher computational overhead, making our approach much faster than those methods.
Table 4: Cost comparison on the mTSP with different numbers of salesmen $M$ and number of nodes NN.
NN $50$ $100$ $200$ Gap(%)
$M$ 5 7 10 5 10 15 10 15 20
LKH3 (solver) 2.00 1.95 1.91 2.20 1.97 1.98 2.04 2.00 1.97 -
OR-Tools (solver) 2.04 1.96 1.96 2.36 2.29 2.25 2.57 2.59 2.59 14.42
GNN-DisPN (g.) 2.14 2.10 1.99 2.56 2.22 2.04 2.97 2.30 2.15 13.45
DAN (g.) 2.29 2.11 2.03 2.72 2.17 2.09 2.40 2.20 2.15 11.75
SchedNet (g.) 2.17 2.07 1.98 2.59 2.13 2.07 2.45 2.24 2.17 10.16
PARCO (g.) 2.12 2.00 1.92 2.47 2.02 1.98 2.28 2.06 1.99 4.43
DAN (s.) 2.12 1.99 1.95 2.55 2.05 2.00 2.29 2.13 2.07 6.13
SchedNet (s.) 2.07 1.99 1.92 2.43 2.03 1.99 2.25 2.08 2.05 4.29
PARCO (s.) 2.07 1.98 1.91 2.38 1.99 1.98 2.22 2.03 1.98 2.80
Table 5: Results for the mTSPLib. CPLEX results with * are optimal solutions. Otherwise, the best-known upper bound of CPLEX results are reported.
instance_N\_N eil51 berlin52 eil76 rat99 Gap $(\%)$
$M$ 2 3 5 7 2 3 5 7 2 3 5 7 2 3 5 7
[…]
We note that these experiments refer to a previous version of PARCO in which the conflict handler was not properly working and was similar to the random handler – we would expect retraining PARCO on the newest implementation would perform even better. As the scope of this comparison is to compare against decentralized methods, we do not include more recent approaches as ET [84] or DPN [104] which would outperform this old PARCO version – albeit with slower decoding. We expect that recent follow-ups of our work, which allow for stopping actions by adding special tokens to the embedding, as MACSIM [68], would outperform ET and DPN, perhaps considerably. We leave this as an interesting direction for future work.
C.3 Convergence Rates
Table 6: Convergence rates in different problems: cost as a percentage of training budget.
10% 25% 50% 75% 100%
HCVRP 5.13 $\pm$ 0.09 5.00 $\pm$ 0.06 4.90 $\pm$ 0.05 4.88 $\pm$ 0.03 4.79 $\pm$ 0.03
OMDCPDP 46.23 $\pm$ 0.21 45.57 $\pm$ 0.15 45.34 $\pm$ 0.11 44.87 $\pm$ 0.10 44.48 $\pm$ 0.09
FFSP 95.45 $\pm$ 0.52 94.32 $\pm$ 0.31 92.88 $\pm$ 0.20 92.15 $\pm$ 0.12 91.48 $\pm$ 0.08
Table˜6 shows the convergence rate of PARCO across different problems on validation datasets at different training budgets (in percentage). PARCO is robust during training and converges stably.
C.4 XXL Instances
To further strengthen our results, we have tested PARCO for even larger scales in both the number of nodes and the number of agents. We generated 16 new instances for $N=5000$ nodes and 3 different values of $M$ agents (a total of 48 instances) of OMDCPDP. We evaluated OR-Tools with 1 hour of runtime and greedy performance for HAM and PARCO.
Table 7: Large-scale generalization results for OMDCPDP with $N=5000$.
$M=500$ $M=750$ $M=1,000$
Obj. Gap Time Obj. Gap Time Obj. Gap Time
OR-Tools 5575.73 134.06% 3600s 5127.46 115.24% 3600s 4974.81 188.10% 3600s
HAM 4813.99 102.08% 17.4s 3732.06 97.33% 19.5s 3258.26 88.69% 22.3s
PARCO 2382.22 0.0% 0.21s 1891.28 0.0% 0.21s 1726.78 0.0% 0.22s
As shown in Table˜7, PARCO excels at generalization at XXL scales with $50\times$ the number of nodes and agents seen during training and up to $1,000$ agents. Thanks to its massively parallel structure, PARCO can solve such instances in a fraction of a second with better results than OR-Tools – at a $10,000\times$ speedup. This makes PARCO ideal for real-world, real-time, large-scale complex problems.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims on the abstract and introduction accurately reflect the paper’s contribution and scope (see methodology of §˜4 and experimental results §˜5).
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We include a paragraph in §˜6 discussing limitations and future works.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: This paper does not include theoretical results.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We give a detailed description of our method in §˜4 and discuss the experimental settings in §˜5 with further information in the Appendix. The source code of PARCO is also publicly available.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: The code of PARCO is available at https://github.com/ai4co/parco. We provide detailed instructions on how to reproduce the main experimental results.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We describe the settings of the training and test details in §˜5, supplemented by further details for each experiment in Appendix˜B. The original configurations are additionally available in the publicly available code to foster reproducibility and openness.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We report the standard deviation in Fig.˜3 with experiments run on 3 different seeds. For the main results including Table˜1, we adhere to the established standard in machine learning for combinatorial optimization papers, which involves reporting averaged results for several dataset instances each to ensure statistical significance: 1,280, 1,000, and 100 for HCVRP, OMDCPDP, and FFSP, respectively.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We report our resources in §˜5.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: Our research conforms to the NeurIPS Code of Ethics
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. Justification: The adoption of PARCO is not expected to have any negative societal impact. On the contrary, we envision machine learning approaches as PARCO may become more prevalent in solving CO problems and thus lead to positive societal benefits, including reduced resource consumption, better management of logistics such as disaster relief efforts, and democratization of optimization solvers.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: There is no high risk of misuse for our released assets.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: Creators of assets (e.g., code, data, models) used in the paper are properly credited in Table˜3.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: All introduced assets are well documented. We provide clear instructions, including checkpoints release through HuggingFace.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [N/A]
  64. Justification: LLMs are used only for writing, editing, or formatting purposes and do not impact the core methodology, scientific rigor, or originality of the research.